OpenStudy (anonymous):

Given f(x) = 3x^2 +x - 1 and g(x) = 4x+ 5, find (gf)(x).

OpenStudy (anonymous):

(gf)(x)=g(f(x))

OpenStudy (anonymous):

so would that be 4x+5(3x^2+x-1) ?

OpenStudy (anonymous):

\[ g(x) = 4 x + 5\\ g(f(x))= 4 f(x) + 5 \] Can you finish it?

OpenStudy (anonymous):

Oh, wo. okay. so its 12x^2+4x+1 ?

OpenStudy (anonymous):

That is right

OpenStudy (anonymous):

hmm, that isn't one of my answer choices.. I think it's supposed to be like this. 4x(3x^2 +x - 1)+ 5 = *12x^3 + 4x^2 -4x + 5*

OpenStudy (anonymous):

a. 3x^2 + 5x + 4 b. 12x^3 + 4x^2 - 4x c. 12x^3 + 19x^2 + x - 5 d. 12x^3 + 19x^2 + 9x - 5 those are the options it gives me.

OpenStudy (anonymous):

this is what its asking (3x^2 + x - 1)(4x + 5) can you solve that?

OpenStudy (anonymous):

Oh, well yeah! that's easy 12x^3 +19x^2 + x - 5

OpenStudy (anonymous):

You were asked to solve g(x) f(x).

OpenStudy (anonymous):

there ya go! :) just name g as the one problem and f as the other, if it says (g + f)(x) then you need to add the two equations. if it says g(f(x)) THEN you would substitute f for x in g's equation

OpenStudy (anonymous):

so if i was to solve g(x) f(x) it would be this: (4x + 5)(3x^2 + x - 1) @eliassaab

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

So that is 12x^3 +19x^2 + x - 5

OpenStudy (anonymous):

\[ (4 x+5) \left(3 x^2+x-1\right)=12 x^3+19 x^2+x-5 \]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ahh you mean (g. f)(x)